OpenStudy (anonymous):

How many liters are in 100 milliliters of water?

OpenStudy (anonymous):

There are 0.1 liters in 100milliliters of water because 1L = 1000mL

OpenStudy (ghazi):

1L = 1000ml or 1000ml=1l so 1 ml = 1/1000 l therefore 100 ml = \[\frac{ 1 }{ 1000 }*100=\frac{ 1 }{ 10 }=0.1L\]